OpenStudy (anonymous):

I need help im helping my kids to finish with this, i need this fast im in a hury! thanks 1. n/4-1=6 2. 6m-1+2m=15 3. 2(7x-6)=16 4. 8(c + 3) –4c = – 5. Apex car rental charges $25 a day plus $.15 per mile for a compact car. Tessie owes $61.40 for a two-day rental. Which equation could be used to find the number of miles she drove? 25m + .15 = 61.40 25m + 2(.15) = 61.40 25 + .15m = 61.40 25(2) + .15m = 61.40 A cell phone company charges $40 a month plus $.10 a minute for any minutes over 200. Which equation shows m, the number of minutes, if the cost for December was $43.10?

OpenStudy (jtvatsim):

1. n/4 - 1 = 6 Add 1 to both sides: n/4 = 7. Multiply by 4 on both sides: n = 28. 2. 6m-1+2m=15 Add 6m and 2m together: 8m - 1 =15. Add one to both sides: 8m = 16 Divide by 8 on both sides: m = 2. 3. 2(7x-6)=16 Divide by 2 on both sides: (7x-6)=8. Remove parenthesis: 7x-6 =8. Add 6 to both sides: 7x = 14. Divide by 7 on both sides: x = 2.

OpenStudy (jtvatsim):

missing info on 8(c + 3) –4c = –(need to put something here)???

OpenStudy (jtvatsim):

5. We know that Tessie used the car for 2 days so she spent at least $50 (that's $25 per day for 2 days). Then she gets charged $.15 per mile driven. That's .15m (since 1 mile would cost .15, 2 miles would cost .30 = 2(.15), etc.) So the equation will be $50 plus an extra .15m equals the total cost $61.40. That is, 25(2) + .15m = 61.40.

OpenStudy (jtvatsim):

Last question. The cost for December was 43.10, so the total charged for minutes is only 3.10. That's 43.10 - 40 (the charge per month). So, the total number of minutes over 200 is found by .10n = 3.10, which solves to n = 31. So the number of minutes used is 200 + 31 = 231.